# SpecExec: Massively Parallel Speculative Decoding for Interactive LLM Inference on Consumer Devices

### v2

@@ Appendix C Additional Implementation Details
Our system design follows the following loop: i) Load the draft model onto GPU memory and generate a draft tree, ii) Load the main model (several layers at a time, if using offloading) to compute probabilities for the draft tree tokens, iii) choose accepted tokens following the verification procedure.
@@
When running the main model, we process all draft tokens in parallel by constructing a merged attention mask, similar to [5]. We prefetch the first few layers of the main model during speculation to speed up the procedure. We also load subsequent LLM layers in parallel, while the previous layers compute their activations, as described in [16, 19].
@@
Finally, we keep the past KV caches of both draft and main models in GPU memory at all times. We chose this because most modern language models use Group Query Attention, making caches relatively small for short prompts. When dealing with longer prompts or smaller GPU memory, one can reduce memory usage by offloading these KV caches into RAM. The draft model caches are only needed on GPU during the first stage when generating the draft tokens. In turn, the main model caches can be loaded alongside their transformer layers.
@@
The practical implementation of this algorithm is slightly different depending on the hardware configuration. Running SpecExec on a system with GPU with RAM offloading works best with relatively fewer draft tokens, while longer offloading (to SSD or when using float16 precision weights) works best with larger token budgets.
@@
As for the quantization scheme, while there are better quantization algorithms [30, 29, 25], we chose GPTQ since it is popular among practitioners. Still, we believe that our experimental results will generalize to other quantization algorithms. In addition to the main model, we also quantize the draft (7B) model using the same GPTQ algorithm. The optimal choices of the quantization methods will vary as new methods or faster implementations come around.
@@
The offloading experiments require sufficient RAM to hold whole model. In case of Llama-2-70B in 16 bit this is at least 140 GB, but in practice 192GB would be recommended to fit the draft model, caches and OS. Our code is based on industry standard PyTorch and Transformers libraries.
Appendix D Ablation: Acceptance with Different Draft Models
@@
In Section 5.2, we evaluate SpecExec and SpecInfer with 7B draft models based on observations about their coverage probabilities. Here, we further compare these models in terms of the number of accepted tokens for different SpecExec batch sizes. We report this comparison in Figure 5 using the same OpenAssistant dataset as in the main experiments using the recommended temperature (0.6) and nucleus size (0.9). Similarly to Figure 2, the 7B model significantly outperforms both JackFram/llama-160m and TinyLlama 1.1B Chat. This is true both for the original 7B model and the one quantized to 4 bits with GPTQ. Curiously, the full unquantized 13B model still scores slightly more accepted tokens, though at the cost of 26GB memory footprint that is inaccessible to modern consumer GPUs.
@@
Figure 5: Number of accepted tokens as a function of draft size $B$ for Llama 2 70B chat target model with a selection of draft models.
Appendix E On The Suboptimality of Beam Search
@@
In our preliminary experiments, we tried to construct the optimal draft tree using top-k beam search decoding [43]. However, we observed that the algorithm performed significantly worse than expected and often plateaued as we increased the maximum beam search length. Here, we describe the analysis of this problem that eventually led us to Algorithm 2.
@@
Figure 6: The average number of accepted tokens per speculation and verification phases as a function of beam size and maximum length. Measured on OpenAssistant conversations and WikiText-2 articles (Right) for running with recommended generation parameters (temperature 0.6, top-p 0.9). (Left) standard beam search decoding, (Right) beam search without pruning out-of-beam tokens
@@
Figure 6 (left) reports a grid where each cell is the number of accepted tokens for a version of SpecExec that uses beam search instead of parallel SSSP. The horizontal grid axis corresponds to beam size (aka the number of beams) and the vertical axis depicts max length. The left grid shows standard beam search decoding that returns beam size most likely sequences. In turn, the right grid uses a modified search algorithm that starts the same way as beam search but does not prune any partial hypotheses that did not make it into the final beam.
@@
As we can see, standard beam search decoding is sub-optimal for SpecExec in the sense that it can be outperformed with trivial modifications. In turn, Algorithm 2 is a generalization of the version on the right that does not need to be manually tuned for length and width, but instead expands the graph optimally to maximize the coverage probability.
Appendix F Application to in-memory inference
@@
While this was not the primary focus area of our research, the SpecExec model can also deliver competitive speedup in inference without offloading. While these speedups are less impressive than those for offload settings, they are still competitive when compared to recent works such as [7].
@@
SL-1.3B / Vicuna-33b oasst 0.6 SX 128 5.33 31.6 2.15x
oasst 0 SX 128 5.4 32.94 2.24x
@@
wikitext 0.6 SX 128 4.87 30.19 1.90x
wikitext 0 SX 128 5.24 33.15 2.08x
Appendix G Effect of draft size on generation speed
Here, we explore the relationship between speed and draft tree size. A larger draft tree size allows for a greater number of tokens to be generated per step (see Figure 7 (left)). However, beyond a certain size threshold (hundreds or low thousands of tokens, depending on the model and GPU), the time required for generation increases at an accelerating rate (see Figure 1 (right)). Consequently, the optimal draft tree size is typically smaller than the size that maximizes the token acceptance rate. In Figure 7 (right), the optimal draft tree size is 128–512 for SpecInfer and 1024–2048 for SpecExec.
Figure 7: Acceptance counts vs draft size (left), speed vs draft size (right). Llama 2-7B draft model, offloaded Llama 2-70B target model, MTBench dataset, t=0.6 and top-p=0.9.

### v3

@@ Appendix D Additional Implementation Details
    Load the main model (several layers at a time, if using offloading) to compute probabilities for the draft tree tokens;
  3. 3.
    Choose the accepted tokens following the verification procedure.
@@
When running the main model, we process all draft tokens in parallel by constructing a merged attention mask, similar to Miao et al. [2023]. We prefetch the first few layers of the main model during speculation to speed up the procedure. We also load subsequent LLM layers in parallel, while the previous layers compute their activations, as described in Pudipeddi et al. [2020], Aminabadi et al. [2022].
@@
Finally, we keep the past key/value caches of both draft and main models in GPU memory at all times. We chose this because most modern language models use grouped-query attention Ainslie et al. [2023], making caches relatively small for short prompts. When dealing with longer prompts or smaller GPU memory, one can reduce memory usage by offloading these KV caches into RAM. The draft model caches are only needed on GPU during the first stage when generating the draft tokens. In turn, the main model caches can be loaded alongside their transformer layers.
@@
The optimal implementation of this algorithm is slightly different depending on the hardware configuration. Running SpecExec on a system with GPU with RAM offloading works best with relatively fewer draft tokens, while longer offloading (to SSD or when using float16 precision weights) works best with larger token budgets.
@@
As for the quantization scheme, while there are better quantization algorithms Lin et al. [2023], Dettmers et al. [2023], Chee et al. [2023], we chose GPTQ since it is popular among practitioners. Still, we believe that our experimental results will generalize to other quantization algorithms. In addition to the main model, we also quantize the draft (7B) model using the same GPTQ algorithm. The optimal choices of the quantization methods will vary as new methods or faster implementations appear.
@@
The offloading experiments require sufficient RAM to hold whole model. In case of Llama-2-70B in 16 bit, this is at least 140 GB, but in practice 192 GB would be recommended to fit the draft model, caches and memory of other processes. Our code is based on industry standard PyTorch Paszke et al. [2019] and Transformers Wolf et al. [2019] libraries.
Appendix E Ablation: Acceptance with Different Draft Models
@@
In Section 5.2, we evaluate SpecExec and SpecInfer with 7B draft models based on the observations about their coverage probabilities. Here, we further compare these models in terms of the number of accepted tokens for different SpecExec batch sizes. We report the results of this comparison in Figure 7 using the same OpenAssistant dataset as in the main experiments using the recommended temperature (0.6) and nucleus size (0.9). Similarly to Figure 2, the 7B model significantly outperforms both JackFram/llama-160m and TinyLlama 1.1B Chat. This is true both for the original 7B model and the one quantized to 4 bits with GPTQ. Curiously, the full unquantized 13B model still obtains slightly more accepted tokens, though at the cost of 26GB memory footprint that is inaccessible to modern consumer GPUs.
@@
Figure 7: Number of accepted tokens as a function of the draft size $B$ for the Llama 2-70B Chat target model and different draft models.
Appendix F On The Suboptimality of Beam Search
@@
In our preliminary experiments, we tried to construct the optimal draft tree using top-k beam search decoding Graves [2012]. However, we observed that the algorithm performed significantly worse than expected and often plateaued as we increased the maximum beam search length. Here, we describe the analysis of this problem that eventually led us to Algorithm 2.
@@
Figure 8: The average number of accepted tokens per speculation and verification phases as a function of beam size and maximum length. The measurements are obtained on OpenAssistant conversations (Left) and WikiText-2 articles (Right) for running with recommended generation parameters (temperature 0.6, top-p 0.9). (Left) standard beam search decoding, (Right) beam search without pruning out-of-beam tokens.
@@
Figure 8 (left) reports a grid where each cell is the number of accepted tokens for a version of SpecExec that uses beam search instead of parallel SSSP. The horizontal grid axis corresponds to beam size (also known as the number of beams), and the vertical axis depicts maximum length within a beam. The left grid shows standard beam search decoding that returns beam size most likely sequences. In turn, the right grid uses a modified search algorithm that starts the same way as beam search but does not prune any partial hypotheses that did not make it into the final beam.
@@
As we can see, standard beam search decoding is suboptimal for SpecExec in the sense that it can be outperformed with trivial modifications. In turn, Algorithm 2 is a generalization of the version on the right that does not need to be manually tuned for length and width, but instead expands the graph optimally to maximize the coverage probability.
Appendix G Application to in-memory inference
@@
While this was not the primary focus area of our research, the SpecExec method can also deliver measurable speedups in inference without offloading. While these speedups are less impressive than those for offload settings, they are still competitive when compared to recent works such as Chen et al. [2024a].
@@
SL-1.3B / Vicuna-33B OASST-1 0.6 SX 128 5.33 31.6 2.15x
OASST-1 0 SX 128 5.4 32.94 2.24x
@@
WikiText-2 0.6 SX 128 4.87 30.19 1.90x
WikiText-2 0 SX 128 5.24 33.15 2.08x
Appendix H Drafting penalty effects
To verify the method’s robustness, we ran a series of experiments with penalties excluding the use of specific tokens. The same penalty scheme was applied to both draft and target models, and the expectation is that the models should be able to run SpecExec effectively. To verify this claim, we ran a series of experiments with penalties excluding use of fewer or more tokens. For these experiments, we penalized all tokens that start from the letter “r” (left) or all tokens that contain the letter “r” (right). Here we used the Llama 2-7B Chat target model with TinyLlama-1.1B Chat draft model, t=0.6, p=0.9, MT-Bench dataset.
The results of these experiments can be found in Figure  9. We found that our method’s performance (measured in terms of accepted tokens per iteration) stays stable only with lightweight penalties, yet heavier penalties reduce the absolute speedups. Looking at the generated samples, we observed that while with lighter penalties, the model is able to work around the restrictions and generate reasonable text, with heavier penalties the quality deteriorated as the model skipped or replaced tokens seemingly at random. Stronger penalties affect the quality of the generated text and naturally make the task harder for the draft model. Thus, we attribute the lower performance with a heavy penalty to this perplexity increase rather than to the penalty directly.
Figure 9: Acceptance rate in generation with token penalty: “don’t start words with “R”” (left) and “don’t use the letter “R”” (right); Llama 2 7B Chat (+ TinyLlama-1.1B Chat draft) model (t=0.6, p=0.9), MT-Bench dataset. The rest of the experimental configuration is the same as in Figure 1.